OpenStudy (anonymous):

Assume womens heights are normally distributed with a mean of 63.4inches and a standard deviation of 2.3inches. if 1 woman is randomly selected what is the probability that her height is less than 64in?

OpenStudy (kirbykirby):

If \(X\) is women's height in inches, then the problem is asking for \(P(X<64)\). Since \(X\) is normally distributed, you can standardize it and then you a standard normal table to find the probability: \[P(X<64)= P\left( \frac{X-63.4}{2.3}<\frac{64-63.4}{2.3}\right)=P(Z<0.26)\]